# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 10                  $\pi=(\pi_{1},\cdots,\pi_{i-1},\pi_{i+1},\cdots,\pi_{n},\pi_{i})$;
In the analysis of results about submodular binary valuation functions, we skillfully use submodularity and the properties of the matroid. Then, we introduce the relationship of submodularity and matroid: a function is submodular binary if and only if it is a rank function of a matroid.
@@
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, where $v_{i}$ is submodular and binary for all $i\in N$. Fix an arbitrary round $k$, let $T=\{e_{1},\cdots,e_{k}\}$ be the set of items that have already arrived, denoted by ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ and $A^{k}_{0}$ the allocation and the set of all unallocated items at the end of round $k$ respectively. In the algorithm procedure, the arriving item is allocated to an agent who thinks that the marginal utility of this item is $1$ if such an agent exists. By Claim 1, $A^{k}_{i}$ is non-wasteful. For each submodular binary valuation function $v_{i}$, there exists a matroid such that $v_{i}$ is the rank function of this matroid. let $\mathcal{M}_{i}=(T,\mathcal{F}_{i})$ be the corresponding matroid of $v_{i}$. We have $v_{i}(A^{k}_{i})=|A^{k}_{i}|$, and $A^{k}_{i}\in\mathcal{F}_{i}$.
@@
Firstly, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-EF1. Consider any two agents $i,j\in N$, $i\neq j$. If $v_{i}(A^{k}_{j})>v_{i}(A^{k}_{i})$, by the property (3) of the matroid, there must exist an item $e\in A^{k}_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$. Let $B_{j}=\{e:e\in A^{k}_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$ and $B_{j}^{*}\in\arg\max\{|S|:S\subseteq B_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$. We will show $|B_{j}|\leq|A^{k}_{i}|+1$. Assume that $|B_{j}|\geq|A^{k}_{i}|+2$, $j$ has priority higher than $i$ at least $|B_{j}|\geq|A^{k}_{i}|+2$ times, which contradicts the order in the algorithm. Combining the property (3) of the matroid, we have $v_{i}(A^{k}_{j})=|A^{k}_{i}|+|B_{j}^{*}|\leq|A^{k}_{i}|+|B_{j}|\leq 2|A^{k}_{i}|+1$. Thus, ${A^{k}}$ is $\frac{1}{2}$-EF1.
@@
Secondly, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-MMS. Consider a MMS partition $X=(X_{1},\cdots,X_{n})$ for agent $i$ of item set $T$, we assume that the share of agent $i$ in ${A^{k}}$ is less than half of the maximin share of $i$, i.e. $v_{i}(A^{k}_{i})\leq\frac{1}{2}v_{i}(X_{j})$ for any $j\in N$. By the property (3) of the matroid, there exists an item $e\in X_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$. Let $C_{j}=\{e:e\in X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$, $C_{j}^{*}\in\arg\max\{|S|:S\subseteq C_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$. We have that $v_{i}(X_{j})=v_{i}(A^{k}_{i}\cup C_{j}^{*})=v_{i}(A^{k}_{i})+v_{i}(C_{j}^{*})=|A^{k}_{i}|+|C_{j}^{*}|$. Since $v_{i}(A^{k}_{i})=|A^{k}_{i}|\leq\frac{1}{2}v_{i}(X_{j})$, $|C_{j}|\geq|C_{j}^{*}|>|A^{k}_{i}|+1$. Then, $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$. We have $|\cup_{j\in N}C_{j}|=|\{e:e\in\cup_{j\in N}X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\{e:e\in T,A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\cup_{j\in N}B_{j}|\leq(n-1)(|A^{k}_{i}|+1)$, it contradicts the inequality $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$. Thus, ${A^{k}}$ is $\frac{1}{2}$-MMS.
@@
Finally, we show that ${A^{k}}=(A^{k}_{1},\cdots,A^{k}_{n})$ is $\frac{1}{2}$-USW. We make a slight adjustment to the Algorithm 2: when a new item arrives, if the marginal utility is 0 for all agents, then assign the item to agent $1$ without updating the agents’ order. Then, we will show that ${A^{k^{\prime}}}=(A^{k}_{1}\cup A^{k}_{0},\cdots,A^{k}_{n})$ is the output of the adjusted algorithm. We only need to prove that in each round $s\in[k]$, if the marginal utility of the arrived item $e_{s}$ to $A^{s-1}_{1}$ is $1$, $v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1$, by the monotonicity of $v_{i}$, $v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0}\cup\{e_{s}\})\geq v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1=v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0})+1$. Then, the marginal utility to $A^{s-1}_{1}\cup A^{s-1}_{0}$ is also $1$. So, ${A^{k^{\prime}}}=(A^{k}_{1}\cup A^{t}_{0},\cdots,A^{k}_{n})$ is the output of the adjusted algorithm. Now, we will show that ${A^{k^{\prime}}}$ is $\frac{1}{2}$-USW. The remaining proof part is based on the idea from [20]. Consider the first item $e_{1}$, we assume that $e_{1}$ is allocated to agent $i^{*}$ during the execution of the algorithm and $v_{i^{*}}(e_{1})=1$, $A_{i^{*}}^{1}=\{e_{1}\}$ and $A_{i}^{1}=\emptyset$ for $i\neq i^{*}$. Let ${Y}=(Y_{1},\cdots,Y_{n})$ be an allocation achieving the optimal USW of instance $I=(T=\{e_{1},\cdots,e_{k}\},N,(v_{i})_{i\in N})$, and we assume that $e_{1}$ is allocated to agent $i^{o}$ in $Y$, $e_{1}\in Y_{i^{o}}$. Next, we construct a new instance ${I^{(1)}}=(T^{\prime},N,(v_{i}^{(1)})_{i\in N})$, where $v_{i^{*}}^{(1)}={\Delta^{i^{*}}_{e_{1}}}$ and $v_{i}^{(1)}=v_{i}$ for all $i\neq i^{*}$. In $I^{(1)}$, $T^{\prime}=T\setminus\{e_{1}\}$, that is, the first item arriving is $e_{2}$. Input the instance $I^{(1)}$ into the adjusted algorithm. Let initialized order of $N$ be $\pi=(\pi_{1},\cdots,\pi_{n},\pi_{i^{*}})$, the initialized allocation be ${D^{0}}=(D^{0}_{1},\cdots,D^{0}_{n})$ , where $D_{i^{*}}=\{e_{1}\}$ and $D_{i}=\emptyset$ for all $i\neq i^{*}$. Assume ${D^{k}}=(D^{k}_{1},\cdots,D^{k}_{n})$ is the allocation at the end of round $k$. We have ${D^{k}}={A^{k^{\prime}}}$, then $USW({D^{k}})=USW({A^{k^{\prime}}})=USW({A^{k}})$. Now, we costruct an allocation ${S}=(S_{1},\cdots,S_{n})$ for instance $I^{\prime}$ as follows: $S_{i^{o}}=Y_{i^{o}}\setminus\{e_{1}\}$, $S_{i}=Y_{i}$ for all $i\neq{i^{o}}$. It holds $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})=v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$. By the monotonicity of $v_{i^{*}}$, $v_{i^{*}}^{(1)}(S_{i^{*}})=v_{i^{*}}(Y_{i^{*}}\cup\{e_{1}\})-v_{i^{*}}(\{e_{1}\})\geq v_{i^{*}}(Y_{i^{*}})-v_{i^{*}}(\{e_{1}\})$. $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$, where the first inequality holds because of the submodularity and the second inequality holds because of $v_{i^{o}}(\{e_{1}\})\leq v_{i^{*}}(\{e_{1}\})$, Let ${Y^{(1)}}=(Y_{1}^{(1)},\cdots,Y_{n}^{(1)})$ be an allocation achieving the optimal USW of instance $I^{(1)}$. It holds that $USW({Y^{(1)}})\geq USW(S)=\cup_{i\in N}v_{i}^{(1)}(S_{i})=\cup_{i\in N,i\neq i^{*},i\neq i^{o}}v_{i}^{(1)}(S_{i})+v_{i^{o}}^{(1)}(S_{i^{o}})+v_{i^{*}}^{(1)}(S_{i^{*}})\geq USW(Y)-2v_{i^{*}}(\{e_{1}\})$. By lemma 1 in [20], $v_{i^{*}}^{(1)}$ is also submodular valuation. So, by constructing the new instances $I^{(2)},\cdots,I^{(k-1)}$ and inducting each round based on the items that arrive, we have $USW(A^{k})=USW(A^{k^{\prime}})=USW(D^{k})=USW(Y^{k})\geq\frac{1}{2}USW(Y).$
@@
For submodular binary valuation functions, no deterministic online algorithm can achieve $(\frac{1}{2}+\epsilon)$-EF1 or $(\frac{1}{2}+\epsilon)$-MMS, and non-wastefulness simultaneously for any $\epsilon>0$.
@@
Consider an online allocation instance, where there are two agents $\{1,2\}$ with identical valuations $v$, and the online items belong to several categories $C_{1},\cdots,C_{l}$. In Figure 1, categories are distinguished by colours, and edges of different colours represent items of different categories. Denote a partition matroid $\mathcal{M}=(T,\mathcal{F})$, =$\mathcal{F}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~{}k\in[l]\}$. The valuation $v$ is the rank function of the matroid $\mathcal{M}$; for any item set $S$, the value of $S$ is the number of items of the maximal independent set of $S$. For the first arrived item $e_{1}$, the marginal utilities for agent $1$ and agent $2$ are $1$. Without loss of generality, we allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, to avoid breaking fairness, we must allocate it to agent $2$. For the third item $e_{3}$, the marginal utilities for agent $1$ and agent $2$ are $1$. We will discuss the following two cases.
@@
Case 1: Allocate $e_{3}$ to agent $2$, shown in Figure 1(1). We must allocate the fourth item, $e_{4}$, to agent $2$ due to the requirement of non-wastefulness. Then, $A^{4}_{1}=\{e_{1}\}$,$A^{4}_{2}=\{e_{2},e_{3},e_{4}\}$. We have that, $v(A^{4}_{1})=\frac{1}{2}v(A^{4}_{2}\setminus\{e\})$ for any $e\in A^{4}_{2}$, and $v(A^{4}_{1})=\frac{1}{2}\mathsf{MMS}_{1}$.
@@
Case 2: Allocate $e_{3}$ to agent $1$, shown in Figure 1(2); we must allocate the fourth item, $e_{4}$, to agent $1$ due to the requirement of non-wastefulness. Then, $A^{4}_{1}=\{e_{1},e_{3},e_{4}\}$,$A^{4}_{2}=\{e_{2}\}$. We have that, $v(A^{4}_{2})=\frac{1}{2}v(A^{4}_{1}\setminus\{e\})$ for any $e\in A^{4}_{1}$, and $v(A^{4}_{2})=\frac{1}{2}\mathsf{MMS}_{2}$. ∎

### v2

@@ 10       $\pi=(\pi_{1},\ldots,\pi_{i-1},\pi_{i+1},\ldots,\pi_{n},\pi_{i})$;
In the analysis of results about submodular binary valuation functions, we effectively leverage submodularity and the properties of the matroid. There exists a relationship of submodularity and matroid: a function is submodular binary iff it is a rank function of a matroid Schrijver et al. (2003).
@@
Given an online fair allocation instance $I=(T,N,(v_{i})_{i\in N})$, where $v_{i}$ is submodular and binary for all $i\in N$. Fix an arbitrary round $k$, let $T^{k}=\{e_{1},\ldots,e_{k}\}$ be the set of items that have already arrived, denoted by ${\bf A}^{k}=(A^{k}_{1},\ldots,A^{k}_{n})$ and $A^{k}_{0}$ the allocation and the set of all unallocated items at the end of round $k$ respectively.
In Algorithm 2, the arriving item is allocated to an agent who thinks that the marginal value of this item is $1$ if there exists such an agent. By Lemma 1, $\textbf{A}^{k}$ is NW. For each $v_{i}$, there exists a matroid $\mathcal{M}_{i}=(T^{k},\mathcal{F}_{i})$ such that $v_{i}$ is the rank function of $\mathcal{M}_{i}$. We have $v_{i}(A^{k}_{i})=|A^{k}_{i}|$, and $A^{k}_{i}\in\mathcal{F}_{i}$.
@@
First, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-EF1. Consider any two agents $\in N, iji\neq j$. If $v_{i}(A^{k}_{j})>v_{i}(A^{k}_{i})$, by the property (3) of the matroid, there must exist an item $e\in A^{k}_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$. Let $B_{j}=\{e:e\in A^{k}_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$ and $B_{j}^{*}\in\arg\max\{|S|:S\subseteq B_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$. We will show $|B_{j}|\leq|A^{k}_{i}|+1$. Assume that $|B_{j}|\geq|A^{k}_{i}|+2$, $j$ has the priority higher than $i$ at least $|B_{j}|\geq|A^{k}_{i}|+2$ times, which contradicts the order in the algorithm. Combining the property (3) of the matroid, we have $v_{i}(A^{k}_{j})=|A^{k}_{i}|+|B_{j}^{*}|\leq|A^{k}_{i}|+|B_{j}|\leq 2|A^{k}_{i}|+1$. Thus, ${\bf A}^{k}$ is $\frac{1}{2}$-EF1.
@@
Second, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-MMS. Consider a MMS partition ${\bf X}=(X_{1},\ldots,X_{n})$ for agent $i$ of item set $T$, we assume that the share of agent $i$ in ${\bf A}^{k}$ is less than half of the maximin share of $i$, i.e. $v_{i}(A^{k}_{i})\leq\frac{1}{2}v_{i}(X_{j})$ for any $j\in N$. By the property (3) of the matroid, there exists an item $e\in X_{j}$ such that $A^{k}_{i}+e\in\mathcal{F}_{i}$. Let $C_{j}=\{e:e\in X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}$ and $C_{j}^{*}\in\arg\max\{|S|:S\subseteq C_{j},A_{i}^{k}\cup S\in\mathcal{F}_{i}\}$. Then, we have $v_{i}(X_{j})=v_{i}(A^{k}_{i}\cup C_{j}^{*})=v_{i}(A^{k}_{i})+v_{i}(C_{j}^{*})=|A^{k}_{i}|+|C_{j}^{*}|$. Since $v_{i}(A^{k}_{i})=|A^{k}_{i}|\leq\frac{1}{2}v_{i}(X_{j})$ holds, we get $|C_{j}|\geq|C_{j}^{*}|>|A^{k}_{i}|+1$. Then, summing up all agents, we have $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$. At last, we get $|\cup_{j\in N}C_{j}|=|\{e:e\in\cup_{j\in N}X_{j},A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\{e:e\in T,A^{k}_{i}+e\in\mathcal{F}_{i}\}|=|\cup_{j\in N}B_{j}|\leq(n-1)(|A^{k}_{i}|+1)$, which contradicts the inequality $|\cup_{j\in N}C_{j}|>n(|A^{k}_{i}|+1)$. Thus, ${\bf A}^{k}$ is $\frac{1}{2}$-MMS.
@@
Finally, we show that ${\bf A}^{k}$ is $\frac{1}{2}$-USW. We make a slight adjustment to the Algorithm 2: when a new item arrives, if the marginal utility is 0 for all agents, then assign the item to agent $1$ without updating the agents’ order. Then, we will show that ${\bf A}^{k^{\prime}}=(A^{k}_{1}\cup A^{k}_{0},\ldots,A^{k}_{n})$ is the output of the adjusted algorithm. We only need to prove that in each round $s\in[k]$, if the marginal value of the arrived item $e_{s}$ to $A^{s-1}_{1}$ is 11, we have $v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1$. By the monotonicity of $v_{i}$, we get $v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0}\cup\{e_{s}\})\geq v_{i}(A^{s-1}_{1}\cup\{e_{s}\})=v_{i}(A^{s-1}_{1})+1=v_{i}(A^{s-1}_{1}\cup A^{s-1}_{0})+1$, implying that the item $e_{s}$’s marginal value to $A^{s-1}_{1}\cup A^{s-1}_{0}$ is $1$. Thus, ${\bf A}^{k^{\prime}}=(A^{k}_{1}\cup A^{t}_{0},\ldots,A^{k}_{n})$ is the output of the adjusted algorithm. Since $v_{1}(A^{k}_{1}\cup A^{t}_{0})=v_{1}(A^{k}_{1})$, we have $USW({\bf A}^{k})=USW({\bf A}^{k^{\prime}})$. In other words, we only need to show that ${\bf A}^{k^{\prime}}$ is $\frac{1}{2}$-max-USW. The remaining proof part is based on the idea from Lehmann et al. (2001). Consider the first item $e_{1}$, we assume that $e_{1}$ is allocated to agent $i^{*}$ during the execution of the algorithm and $v_{i^{*}}(e_{1})=1$, $A_{i^{*}}^{1}=\{e_{1}\}$ and $A_{i}^{1}=\emptyset$ for $i\neq i^{*}$. Let ${\bf Y}=(Y_{1},\ldots,Y_{n})$ be an allocation achieving the optimal USW of instance $I=(T^{k},N,(v_{i})_{i\in N})$, and we assume that $e_{1}$ is allocated to agent $i^{o}$ in $Y$, $e_{1}\in Y_{i^{o}}$. Next, we construct a new instance ${I^{(1)}}=(T^{\prime},N,(v_{i}^{(1)})_{i\in N})$, where $v_{i^{*}}^{(1)}={\Delta^{i^{*}}_{e_{1}}}$ and $v_{i}^{(1)}=v_{i}$ for all $i\neq i^{*}$, $T^{\prime}=T^{k}\setminus\{e_{1}\}$, that is, the first item arriving is $e_{2}$, and then input the instance $I^{(1)}$ into the adjusted algorithm. Let initialized order of $N$ be $\pi=(\pi_{1},\ldots,\pi_{n},\pi_{i^{*}})$, the initialized allocation be ${\bf D}^{0}=(D^{0}_{1},\ldots,D^{0}_{n})$ , where $D_{i^{*}}=\{e_{1}\}$ and $D_{i}=\emptyset$ for all $i\neq i^{*}$. Assume that ${\bf D}^{k}=(D^{k}_{1},\ldots,D^{k}_{n})$ is the allocation at the end of round $k$. We have ${\bf D}^{k}={\bf A}^{k^{\prime}}$, then $USW({\bf D}^{k})=USW({\bf A}^{k^{\prime}})=USW({\bf A}^{k})$. Now, we construct an allocation ${\bf S}=(S_{1},\ldots,S_{n})$ for instance $I^{\prime}$ as follows: $S_{i^{o}}=Y_{i^{o}}\setminus\{e_{1}\}$, $S_{i}=Y_{i}$ for all $i\neq{i^{o}}$. It holds that $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})=v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$. By the monotonicity of $v_{i^{*}}$, we have $v_{i^{*}}^{(1)}(S_{i^{*}})=v_{i^{*}}(Y_{i^{*}}\cup\{e_{1}\})-v_{i^{*}}(\{e_{1}\})\geq v_{i^{*}}(Y_{i^{*}})-v_{i^{*}}(\{e_{1}\})$. $v_{i^{o}}^{(1)}(S_{i^{o}})=v_{i^{o}}(Y_{i^{o}}\setminus\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{o}}(\{e_{1}\})\geq v_{i^{o}}(Y_{i^{o}})-v_{i^{*}}(\{e_{1}\})$, where the first inequality follows from the submodularity and the second inequality follows from $v_{i^{o}}(\{e_{1}\})\leq v_{i^{*}}(\{e_{1}\})$. Let ${\bf Y}^{(1)}=(Y_{1}^{(1)},\ldots,Y_{n}^{(1)})$ be an allocation achieving the optimal USW of instance $I^{(1)}$. It holds that $USW({\bf Y}^{(1)})\geq USW({\bf S})=\cup_{i\in N}v_{i}^{(1)}(S_{i})=\cup_{i\in N,i\neq i^{*},i\neq i^{o}}v_{i}^{(1)}(S_{i})+v_{i^{o}}^{(1)}(S_{i^{o}})+v_{i^{*}}^{(1)}(S_{i^{*}})\geq USW({\bf Y})-2v_{i^{*}}(\{e_{1}\})$. By lemma 1 in Lehmann et al. (2001), $v_{i^{*}}^{(1)}$ is also submodular. So, by constructing the new instances $I^{(2)},\ldots,I^{(k-1)}$ and inducting each round based on the items that arrive, we have $USW({\bf A}^{k})=USW({\bf A}^{k^{\prime}})=USW({\bf D}^{k})=USW({\bf Y}^{k})\geq\frac{1}{2}USW({\bf Y}).$
@@
For any $\epsilon>0$, with submodular binary valuation functions, no deterministic online algorithm can achieve
  • $(\frac{1}{2}+\epsilon)$-EF1 and non-wastefulness,
  • or $(\frac{1}{2}+\epsilon)$-MMS and non-wastefulness,
  • or $(\frac{1}{2}+\epsilon)$-max-USW and non-wastefulness.
@@
First, let us consider EF1 or MMS. Consider an online allocation instance, where there are two agents $\{1,2\}$ with identical valuations $v$, and the online items belong to several categories $C_{1},\ldots,C_{l}$. In Figure 1, categories are distinguished by colours, and edges of different colours represent items of different categories. Denote a partition matroid $\mathcal{M}=(T,\mathcal{F})$, $\mathcal{F}=\{S:S\subseteq T,|S\cap C_{k}|\leq 1,\text{for all}~k\in[l]\}$. The valuation $v$ is the rank function of the matroid $\mathcal{M}$; for any item set $S$, the value of $S$ is the number of items of the maximal independent set of $S$. For the first arrived item $e_{1}$, the marginal values for agent $1$ and agent $2$ are $1$. Without loss of generality, we allocate item $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, in order not to break fairness, we must allocate it to agent $2$. For the third item $e_{3}$, the marginal values for agent $1$ and agent $2$ are $1$. We will discuss the following two cases.
@@
Case 1: Allocate $e_{3}$ to agent $2$, shown in the left graph in Figure 1 Then, $A^{4}_{1}=\{e_{1}\}$$A^{4}_{2}=\{e_{2},e_{3},e_{4}\}$. We have that, $v(A^{4}_{1})=\frac{1}{2}v(A^{4}_{2}\setminus\{e\})$ for any $e\in A^{4}_{2}$, and $v(A^{4}_{1})=\frac{1}{2}\mathsf{MMS}_{1}$.
@@
Case 2: Allocate $e_{3}$ to agent $1$, shown in the right graph in Figure 1; we must allocate the fourth item, $e_{4}$, to agent $1$ due to the requirement of non-wastefulness. Then, $A^{4}_{1}=\{e_{1},e_{3},e_{4}\}$$A^{4}_{2}=\{e_{2}\}$. We have that, $v(A^{4}_{2})=\frac{1}{2}v(A^{4}_{1}\setminus\{e\})$ for any $e\in A^{4}_{1}$, and $v(A^{4}_{2})=\frac{1}{2}\mathsf{MMS}_{2}$.
Next, let us consider USW. Consider an online allocation instance, where there are two agents $\{1,2\}$. For the first arrived item $e_{1}$, the marginal values for agents $1$ and $2$ are both $1$, $v_{1}(\{e_{1}\})=v_{2}(\{e_{1}\})=1$. Without loss of generality, we allocate $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, $v_{1}(\{e_{2}\})=v_{1}(\{e_{1},e_{2}\})=1$$v_{2}(\{e_{2}\})=0$, the marginal values for agents $1$ and $2$ are both 0. No matter which agent gets $e_{2}$, $USW(A^{2})=1$. However, the max-USW is $2$ by allocating $e_{1}$ to agent $2$ and $e_{1}$ to agent $1$. ∎
For an independent interest, we find that there is a side result of the Marginal-Greedy Algorithm, where it can also work for additive binary valuation functions to compute an NW allocation that is EF1, MMS, and max-USW, which demonstrates the robustness of our algorithm.
3.3 Additive Personalized Bi-valued Valuations
In this part, we consider the additive personalized bi-valued valuation function, which is often considered in the literature, such as Ebadian et al. (2022); Aziz et al. (2023). For any agent $i\in N$, her valuation function $v_{i}$ is additive personalized bi-valued if $v_{i}$ is additive and $v_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$.
It is worth investigating whether an NW allocation satisfying EF1 or MMS is achievable with additive personalized bi-valued valuation functions, as the binary case is a special type of bi-valued function. However, we find that even for two agents with additive personalized bi-valued valuations, an NW allocation satisfying EF1 or MMS cannot be guaranteed.
3.3.1 Two Agents
We investigate the setting of additive personalized bi-valued valuations for two agents and establish the optimal non-trivial result for simultaneously approximating both EF1 and MMS. To achieve this objective, we propose the Adapted Envy-Graph Procedure (Algorithm 3), which requires sophisticated and subtle case analysis due to the uncertainty of future item values.
The Adapted Envy-Graph Procedure (AEGP) consists of three sub-algorithms: the Envy-Graph Procedure (EGP), the Preliminary-Breaking-Cycle (PBC), and the Deep-Breaking-Cycle (DBC), which are detailed in the Appendix. The core structure of our algorithm employs the EGP as the primary procedure. However, the EGP has a critical requirement: the input allocation is EF1 and must not contain any envy cycle.
When an envy cycle is detected, we execute the PBC and DBC algorithms as auxiliary procedures in subsequent rounds to allocate future items more effectively. Once the output allocation from either the PBC or DBC algorithm satisfies the EGP algorithm’s precondition, the EGP algorithm resumes execution. The execution flow of this integrated process is illustrated in Figure 2.
Figure 2: The execution flow of EGP, PBC, and DBC algorithms. The dashed line represents the execution of the corresponding algorithm. The solid line means that when some specific cases occur, we execute another algorithm.
The main challenge lies in addressing envy cycles effectively. This issue is straightforward to resolve in offline settings but becomes significantly more complex in online environments where reallocation is infeasible. To overcome this obstacle, we first introduce the PBC algorithm, which is developed through careful analysis of item values over the next two rounds. Specifically, the PBC algorithm can be executed consecutively for at most two rounds. After these two rounds, the PBC algorithm successfully addresses most cases, achieving EF1 allocation without an envy cycle. However, if the allocation remains be $\frac{1}{2}$-EF1, a more strategic approach is required in subsequent rounds to ensure fairer distribution. To address this limitation, we introduce the DBC algorithm. In the DBC algorithm, we establish allocation rules based on the parity of consecutive execution counts. Unlike the PBC algorithm, the DBC algorithm can be executed consecutively across multiple rounds while maintaining the $\frac{1}{2}$-EF1 approximation guarantee, before its output allocation satisfies the precondition of the EGP. In summary, both the PBC and DBC algorithms establish specific allocation rules for future items when an envy cycle emerges. These rules not only ensure EF1 approximation throughout their execution but also continuously monitor whether the EGP precondition can be satisfied and actively work to achieve it when feasible.
Definition 6 (Envy Graph for Goods).
Given an allocation ${\bf A}^{k}$, the corresponding envy graph is defined as $G=(V,E)$, where the vertex set $V$ corresponds to agent set $N$, and a directed edge $(i,j)\in E$ iff agent $i$ envies agent $j$, i.e., $v_{i}(A_{i}^{k})<v_{i}(A_{j}^{k})$. Additionally, the directed cycle in the envy graph is called an envy cycle.
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive personalized bi-valued valuation functions
Output: An NW approximate EF1 and approximate MMS allocation $\bf A$}
1
2Let ${\bf A}=(\emptyset,\ldots,\emptyset)$;
Initialize $\alpha=1$; // Algorithm Indicator
3
Initialize $(\lambda_{1},\mu_{1})$ and $(\lambda_{2},\mu_{2})$, where $\lambda_{1}=\lambda_{2}=1$ and $\mu_{1}=\mu_{2}=0$; ;
  // $\lambda$: the times of consecutive execution for the PBC or DBC algorithm, and $\mu$: the case label in the PBC or DBC algorithm.
4
5Build an envy graph $G$ for two agents;
6 when item $e\in T$ arrives do
7 if $\alpha=1$ then
8
9    ${\bf A},\alpha\leftarrow\mathsf{EGP}(N,{\bf A},\alpha,e)$;
10    if $\alpha=2$ then
       ${\bf A},\alpha\leftarrow\mathsf{PBC}(N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1}))$ ;
          // Envy cycle will appear.
11
12      end if
13
14 else if $\alpha=2$ then
15    ${\bf A},\alpha,(\lambda_{1},\mu_{1})\leftarrow\mathsf{PBC}(N,{\bf A},\alpha,e,(\lambda_{1},\mu_{1}))$;
16 else if $\alpha=3$ then
17    ${\bf A},\alpha,(\lambda_{2},\mu_{2})\leftarrow\mathsf{DBC}(N,{\bf A},\alpha,e,(\lambda_{2},\mu_{2}))$;
18   end if
19
20  Update the envy graph $G$;
21
return $\bf A$};
Algorithm 3 Adapted Envy-Graph Procedure